OpenStudy (anonymous):

Having trouble finding the volume: Solid whose base is the region bounded by y=x^2 and y = 9. And whose cross sections are semi circles perpendicular to the base and parallel with the x-axis. I know the interval should be between 0 and 3, but what confuses me is the semi-circle cross section part.

OpenStudy (anonymous):

The volume will essentially be the sum of the cross-sections' areas. The area of each cross section is \(\dfrac{1}{2}\pi r^2=\dfrac{1}{2}\pi\left(\dfrac{d}{2}\right)^2=\dfrac{1}{8}\pi d^2\), since they are all semicircles. Each cross-section is parallel to the x-axis, which means the diameter of each cross section should be expressed in terms of y: \[V=\int_a^b A(y)~dy\] Each cross-section's diameter is the distance from one end of the parabola to the other: